HOME ESR EXPERIMENTS LABS etc
NucAR lab laser_lab Fritz-Bosch-Lab X-Ray Lab target Beavertail
  Beavertail, all entries  Not logged in ELOG logo
ID Date Author Category Subject Text Attachments
  1   Wed Jun 15 13:11:38 2022 Guy LeckenbyDAQSignal DiagramAttached is the signal mapping that will be
used for the BEAVERTAIL test.
 Beavertail_Signal_Map.pdf 
  2   Fri Jun 17 15:32:47 2022 Guy LeckenbyElectronicsPreamp SettingsBias Select: In general, higher bias resistors
-> lower electronic noise. However, higher
resistance = larger voltage drop due to leakage
 Preamp_Settings.png 
  3   Mon Jun 20 10:36:56 2022 Guy LeckenbyDAQDAQ Startup CommandsThree terminals/processes are required to
record data on a local machine.
1. log in x86l-131. 
  
  4   Tue Jun 21 02:37:07 2022 Guy Leckenby, Iris Dillmann, Chris GriffinGeneralRun notesBeam: 208Pb(82+), 269.0 MeV/u, ~2*10^7 ions
in ESR
 Pocket-drive.pngecooler-current.png208Pb82_Beam-Lifetime.png 
  5   Wed Jun 22 18:01:18 2022 Guy LeckenbyDetectorsDetector ArrangmentAttached is the detector arrangement for the
3 setups used during the tests.
 Detector_Setups_for_Test.pdf 
  6   Wed Jun 22 19:14:12 2022 Guy LeckenbyDAQFBEX Script ChoicesHere were the f_user.c choices we implemented
for this experiment.
 f_user.c 
  7   Tue Jun 28 17:42:53 2022 Guy LeckenbyDAQCabling ChangesIn the previous signal map, I incorrectly
mapped the DSSD channels. To correctly read
out the DSSD for the new test, we will swap
 DSSD_New_Ch.png 
  8   Fri Jul 29 15:33:08 2022 Guy LeckenbyElectronicsResults of Noise TestingUwe and I were able to get some noise testing
done before the ESR was closed with asbestos
exposure. We used the new flange so didn't
 PLEIADES_Noise_-_coupled_to_ESR.jpgPLEIADES_Noise_-_isolated_from_ESR.jpg 
  9   Mon Nov 7 14:14:38 2022 Chris Griffin, Guy LeckenbyElectronicsMSI-8 module detailsMSI-8 module details.

MMPR1 card configurations:
  
  10   Fri Nov 11 10:27:59 2022 Chris Griffin, Guy LeckenbyDetectorsCurrent leakage currentsAttached are photos of leakage currents for
the current GAGG and BGO setups.
This setup uses the new flange (blue gas
 Housing1-GAGG.jpgHousing2-BGO.jpg 
  11   Tue Nov 15 14:30:25 2022 Chris GriffinElectronicsOld flange noiseLast week we tested both the new and old flanges
with the GAGG housing set up described in
the previous ELog entry.
 oldFlange-MPR.jpg 
  12   Thu Nov 17 13:42:16 2022 Chris GriffinDetectorsDetector config changesI have changed the configuration of Housing
2 (BGO) to account for the detector being
fixed (bond wires) and provide a "plug-and-play"
  
  14   Mon May 29 12:23:48 2023 Guy Leckenby, Chris GriffinElectronicsOld flange noise re-testsWhilst constructing a new flange to replace
the old one, we did some tests again to compare
the performance of new vs old flange.
 26.05.23-MHLab-Old_Flange_Biased_No_Grnd_Con_n-sides.PNG26.05.23-MHLab-OF-B-YGC-n-sides.PNG26.05.23-ESR-OF-B-YGC-n-sides.PNG26.05.23-ESR-OF-B-YGC-p-sides.PNG26.05.23-ESR-NF-B-NGC-n-sides.PNG 
  15   Thu Jun 1 09:39:24 2023 Guy Leckenby, Chris GriffinElectronicsNew new flange testWe tested the new new flange using the same
setup as in ELOG entry 14. The noise level
seemed to be somewhere between the middle
 31.05.23-ESR-NNF-YGC-n-sides.PNG31.05.23-ESR-NNF-NGC-n-sides.PNG31.05.23-ESR-NNF-YGC-n-sides-v2.PNG 
  16   Thu Jun 1 10:42:31 2023 Guy LeckenbyElectronicsSummary of Preamp TestingMSI-8 and MPR-16 preamps have nominal differential
output ranges of +/-1.8V and +/-1.2V respectively
at maximum input (ie 5GeV).
 preamp_tests.png 
  17   Tue Dec 12 12:57:16 2023 JanDAQtest lmd-files with pulserWe took a few lmd-files with different settings
of the febex DAQ to provide data for go4
development, when the DAQ is not running.
  
  18   Mon Mar 24 14:44:14 2025 C Griffin, M Cockshutt, I DillmannGeneralApril 2025 experiment setupFurther PLEIADES tests are to be carried out
during E018 and E203. This will be an important
test of PLEIADES operation in isochronous
  
  19   Fri Mar 28 13:31:53 2025 CG, ID, MC, JGGeneralDual PLEIADES installation at ESRPLEIADES detectors were installed in both
the inside and outside pockets after the
dipole behind the gas jet.
 Signal_mapping-default_table.pngPXL_20250327_132426516.jpgPXL_20250328_095538211.jpgPXL_20250328_095849127.jpg 
  20   Fri Mar 28 20:45:08 2025 CGElectronicsMSI-8 configurationsThe MSI-8 modules were reconfigured to properly
instrument the current setup.
 PXL_20250328_183822721.jpgPXL_20250328_184605620.jpgPXL_20250328_191007196.jpgPXL_20250328_190527927.jpgPXL_20250328_183302403.jpg 
  21   Fri Mar 28 22:45:24 2025 CGDetectorsBias problemsAfter testing the biasing scheme for all detectors,
two problems emerge.
 designI-13_biasResponse.pngMSPAD16A_biasResponse.png 
  22   Sat Mar 29 15:51:46 2025 CG, JG, MC, IDGeneralDAQ setup********************************************
** Setup of DAQ x86l-131 (outside pocket)
**
  
  23   Sat Mar 29 17:19:46 2025 CG, JG, MC, IDGeneralSample data fileData files written to /mnt/raw.data/e0203_e0018.

From MBS computers:
  
  24   Sat Mar 29 17:53:13 2025 CG, ID, MCElectronicsBias remote controlThe MHV modules are set up for remote control.
This can be accessed via the mesycontrol
software running on x86l-128.
  
  25   Mon Mar 31 17:55:16 2025 JGDAQDAQ overviewThe DAQ is running of 4 different nodes/PCs,
that can be accessed by ssh using the user
litv-exp.
  
  26   Mon Mar 31 18:21:37 2025 JGDAQESR scaler mappingThis is the list of scaler channels used at
the VME ESR DAQ on x86l-204:
  
  27   Mon Mar 31 18:35:46 2025 JGDAQlmd files on ESR NASThe files of each MBS node or the timesorter
are written to the ESR NAS by the RFIO server.
The server is running on the virtual machine
  
  28   Tue Apr 1 17:18:10 2025 ID, MC, CJGDetectorsBoth PLEIADES detectors in beamWe started inserting the inside and outside
detector, see
https://elog.gsi.de/esr/E0203_E0018/46
  
  30   Tue Apr 1 22:41:52 2025 CGDetectorsClimbing leakage currentsThe experiment has been running stably during
the afternoon.
DAQ rates have remained constant and the
 highLeakageCurrents.png 
  31   Tue Apr 1 23:00:41 2025 CGGeneralRun stopped and detectors moved backDue to the continued rise in leakage currents,
I have stopped the run, moved the detectors
completely back and turned off the 
  
  32   Thu Apr 3 13:09:28 2025 JGDAQ194Pb MBS settingsThe settings of the MBS daq during the first
data taking with fragments around 194Pb (W.Korten
experiment) was not entirely correct:
  
  33   Thu Apr 3 13:24:31 2025 CGDetectorsBias problem diagnosingDuring an access break to the ESR, I made
a few tests with the bias scheme to see if
we could better understand some of the weird
  
  34   Thu Apr 3 14:53:13 2025 IDDetectorsLink to leakage current Google docPlease enter the leakage currents from the
mesycontrol every ~1h into this Google document:
  
  35   Thu Apr 3 17:37:04 2025 CJG, GL, MCGeneralDetectors in position, taking dataThe inner detector was moved into position.
The device control for the outer detector
is not responding and I am not an expert/don't
  
ELOG V3.1.5-fc6679b